OpenStudy (study312):

Solve -7/3x - 3 = -52.

OpenStudy (study312):

\[\frac{ -7 }{ 3 }x - 3 = -52\]

OpenStudy (heart_offire):

are we solving for x

OpenStudy (study312):

It says solve it so I guess

OpenStudy (amy0799):

-7/3=-49 -7x=-147 x=21

OpenStudy (study312):

what happened to the -3?

OpenStudy (amy0799):

you add it to the right side

OpenStudy (study312):

oh you subtracted -3 from both sides

OpenStudy (study312):

oh ok

OpenStudy (study312):

thx

OpenStudy (study312):

so what did you do with the 3 on the bottom of -7?

OpenStudy (amy0799):

multiply it to the right side

OpenStudy (study312):

ok